The base of a triangle is 3 times the length of the altitude to that base. The area of the triangle is 96 cm^2. Find the length of the base.
Answer by unlockmath(1688)

Hello,
First off, we know the area of a triangle is 1/2*base*height. So with that we can plug in the numbers, in this case we'll use 3x for base and x for the height.
1/2(3x)(x)=96 Rewritten it will be:
1/2(3x^2)=96 Multiply both sides by 2 gives us:
3x^2=192 Divide each side by 3 will be:
x^2=64 Square root both sides is:
x=8 cm This is the height and the base is
24 cm
Make sense?
RJ
Check out a new book at:
www.math-unlock.com